OpenStudy (anonymous):

Find the missing measure for a right circular cone given the following information. Find h if r = 5 and V = 100. a)4 b)12 c)30

jimthompson5910 (jim_thompson5910):

V = (pi*r^2*h)/3 100 = (pi*5^2*h)/3 3*100 = pi*5^2*h 300 = 3.14*25*h Can you take it from here?

OpenStudy (anonymous):

So what am I multiplying? 3.14*25?

jimthompson5910 (jim_thompson5910):

yes

jimthompson5910 (jim_thompson5910):

which gives you what?

OpenStudy (anonymous):

78.5

OpenStudy (anonymous):

Just remember your algebra. How do you solve an equation with 1 variable?

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

so 300 = 3.14*25*h becomes 300 = 78.5*h

OpenStudy (theviper):

\[\LARGE{\color{red}{h=\frac{300}{78.5}}}\]

OpenStudy (anonymous):

How is h , 300/78.5?

OpenStudy (anonymous):

Look at the equation again. How do you solve for h?

jimthompson5910 (jim_thompson5910):

you start with 300 = 78.5*h and you divide both sides by 78.5 to get 300/78.5 = h

OpenStudy (anonymous):

Figured it out, thanks guys!:)

OpenStudy (anonymous):

what is the answer? A, B, or C?